OpenStudy (anonymous):

PROBABILITY QUESTION men that would stop and ask for direction : 42% women that would stop and ask for direction : 61% Population estimation, men : 48.2% women : 51.8% 1) calculate probability driver stops and ask for direction 2) given driver stops to ask for directions, determine probability that the driver was a man.

OpenStudy (anonymous):

Okay so let \(A\) be the event you stop and ask for directions and let \(B\) be the event you're a man.

OpenStudy (anonymous):

"men that would stop and ask for direction : 42%"\[ \Pr(A|B)=0.42 \]

OpenStudy (anonymous):

"women that would stop and ask for direction : 61%"\[ \Pr(A|B^C)=0.61 \]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Population estimation, men : 48.2% women : 51.8% \[ \Pr(B)=0.482\\ \Pr(B^C)=0.518 \] Remember when someone says \[\Huge \text{Women and minorities} \]What they are actually saying is \[\Huge \text{Everyone} \]because men are a minority.

OpenStudy (anonymous):

Okay so finally, you need to remember that: \[ \Pr(A|B) = \frac{\Pr(A\cap B)}{\Pr(B)} \]

OpenStudy (anonymous):

In this case, it helps to realize \[ \Pr(A\cap B)=\Pr(A|B)\times \Pr(B) \]

OpenStudy (anonymous):

"1) calculate probability driver stops and ask for direction" What they want is \(\Pr(A)\). We need to use total probability here: \[ \Pr(A)= \Pr(A\cap B) + \Pr(A\cap B^C) \]

OpenStudy (anonymous):

1.03?

OpenStudy (anonymous):

"2) given driver stops to ask for directions, determine probability that the driver was a man." This asks for \(\Pr(B|A)\). It's going to help to know \(\Pr(A)\) for this one since: \[ \Pr(B|A)=\frac{\Pr(B\cap A)}{\Pr(A)} \]

OpenStudy (anonymous):

but i thought probability cant be more than 1?

OpenStudy (anonymous):

You didn't do it right.

OpenStudy (anonymous):

Show me your calculation.

OpenStudy (anonymous):

0.42 + 0.61

OpenStudy (anonymous):

is that wrong? for the first question

OpenStudy (anonymous):

or do i have to divide with the total sum?

OpenStudy (anonymous):

but if its by % the total sum should be one right? im quite confuse

OpenStudy (anonymous):

Okay what you have is \[ \Pr(A|B)+\Pr(A|B^C) \]You want to find \[ \Pr(A\cap B)+\Pr(A\cap B^C) \]

OpenStudy (anonymous):

Remember that \[ \Pr(A|B) = \frac{\Pr(A\cap B)}{\Pr(B)} \]

OpenStudy (anonymous):

You're confusing conditional probability with joint probability.

OpenStudy (anonymous):

Do I need to spell it out more?

OpenStudy (anonymous):

\[ \Pr(A\cap B) = \Pr(A|B)\times \Pr(B) = (0.42)(0.482) \]

OpenStudy (anonymous):

It's very similar for \(\Pr(A\cap B^C)\)

OpenStudy (anonymous):

ah okay i see

OpenStudy (anonymous):

thank you so much! @wio